OpenStudy (anonymous):

Is (pie-over-64) rational or irrational? Explain your reasoning.

OpenStudy (anonymous):

irrational, b/c pi is not a whole number

OpenStudy (anonymous):

pi?

OpenStudy (anonymous):

\(\Large \frac{\pi}{64}=\frac{irrational}{rational}=irrational \)

OpenStudy (anonymous):

yeah pi...

OpenStudy (anonymous):